OpenStudy (anonymous):

the angles of a quardrilateral are in the ratio 1:2:4:5 find the angles. please show steps.

OpenStudy (anonymous):

The angles sum to 360, divide in 12, then 1/12,2/12,4/12.....

OpenStudy (anonymous):

it can be done like this first add all the numbers 1+2+4+5=12 now each angle 1*360/12 2*360/12 4*360/12 5*360/12 hence the angles are 30,60,120,150

OpenStudy (anonymous):

1+2+4+5=12 360/12= 1 in the ratio So if 1=30 degrees then each number is each angle respectively.